Fred found two dollars. He now has five times as much money as he would have had if he lost two dollars. How many dollars did Fred have originally?
I also need a formula for this problem. Thank you!

Answer by stanbon(75887)

Fred found two dollars. He now has five times as much money as he would have had if he lost two dollars. How many dollars did Fred have originally?
----------------------
Let his original amount be "x".
Equation:
x+2 = 5(x-2)
x+2 = 5x-10
4x = 12
x = 3 dollars
-------------------
Cheers,
Stan H.
---------------